# MaskMoE: Boosting Token-Level Learning via Routing Mask in Mixture-of-Experts

### v1

@@ 3.3 Load Balance Loss
MoE models typically require distributed training, where different experts are deployed across various nodes. To avoid load imbalance, where a minority of expert nodes handle the majority of tokens while the majority of expert nodes spend more time idle, this can impact training efficiency. It is generally desirable for the number of tokens processed by different experts to be roughly equal. To achieve load balancing, a load balancing loss is commonly introduced in the training of MoE models. We follow previous works Huang et al. (2024); Fedus et al. (2022) and adopt a widely used loss.
@@
Where $T$ represents the number of tokens in a mini-batch. $\mathbf{p}^{j}$ denotes the probability distribution of the routing output for the j-th token, while $\mathbf{p}_{i}^{j}$ represents the probability value of the i-th expert. It is noteworthy that for infrequent tokens, which have only one visible expert, the load balancing loss does not apply to them. The balancing loss primarily regulates the routing of frequent tokens. Additionally, since infrequent tokens are assigned to experts through random hashing, the overall load is relatively balanced. Therefore, our experiments find that MaskMoE, compared to SMoE, does not experience a significant decrease in training efficiency.
@@ 4 Expertiments
Table 1: Perplexity results of language modeling.
@@ 4.1 Pre-training Dataset
Following previous works, we use the Pile dataset Gao et al. (2021) for pre-training data. The Pile dataset is a large-scale, publicly available corpus for language model pre-training, containing 22 domains and over 825GB of English text data. For our experiments, we use the well-known LLaMA tokenizer for tokenization, with a vocabulary size of $32k$. We follow  Xie et al. (2023a); Su et al. (2024) to calculate the sampling rate for each domain based on the number of tokens after tokenization. Due to the computational budget, we follow the pretraining settings of  Xie et al. (2023a); Su et al. (2024); Huang et al. (2024); Xiong et al. (2024); Lian et al. (2024), all models are pre-trained with 100B tokens.
Then, to identify infrequent tokens and frequent tokens, we calculate the frequency of each token in the training set and sort them in descending order of frequency. We categorize the top tokens that cover $P$of the dataset as frequent tokens, and the remaining $(1-P)$as infrequent tokens. Where $P$ is a tunable hyperparameter.
@@ 4.2 Experimental Setup
Following  Touvron et al. (2023a, b); Huang et al. (2024), we adopt the LLaMA architecture with 24 transformer layers and a hidden layer dimension of 1024. In line with LLaMA Touvron et al. (2023a), we employ the AdamW  Loshchilov and Hutter (2019) optimizer, and applying a cosine learning rate decay schedule.
For the MoE model, following  Roller et al. (2021), we conduct experiments under both single-layer and multi-layer settings. For the single-layer setting, we replace the final FFN layer with the MoE layer. For the multi-layer MoE models, we follow Gshard Lepikhin et al. (2021), alternating MoE layers every other layer, resulting in a total of $12$ MoE layers in this setting. For single-layer MoE models and dense models, we use a learning rate of $43e-$ 4, while for $12$-layer MoE models, we employ a learning rate of $41e-$ 4 to ensure stable convergence of the models. If not specified otherwise, tokens that account for the top 40% of frequency are considered frequent tokens, while the remaining are considered infrequent tokens, i.e., $P=0.4$. For frequent tokens, the visible expert count $V$ is set to 8, and for infrequent tokens, the visible expert count $V$ is set to 1. All of our implementations are based on the DeepSpeed [^1] https://github.com/microsoft/DeepSpeed library Rajbhandari et al. (2022); Rasley et al. (2020), which offers robust support for MoE-distributed training. By default, we enable the random token selection strategy Kim et al. (2021) implemented in the library to facilitate faster model convergence and better runtime efficiency. Following previous works Roller et al. (2021); Fedus et al. (2022); Dai et al. (2022), unless otherwise specified, our experiments only select the top expert.
4.3 Compared Models
Including MaskMoE, we compare five models for validation experiments. Next, we will briefly introduce five compared models.
  • Dense represents a standard Transformer language model with 468M parameters.
  • Hash Layer Roller et al. (2021) signifies that through a random hash method, each token is assigned a fixed routing expert prior to training. Unless otherwise specified, each MoE layer has 64 experts.
  • MaskMoE achieves a balance between representation diversity in MoE layers and thorough parameter training through the routing masking method we proposed. Similarly, in this section, MaskMoE also employs the shared expert architecture, which is identical to that of Share-MoE.
@@ 4.3 Compared Models
Table 2: Performances of language models on downstream benchmarks. The best score is marked in bold, and the second best is underlined.
@@ 4.4 Main Results
In this section, we first present the model’s perplexity (PPL) on the Pile validation set. Then, following Touvron et al. (2023a); Brown et al. (2020); Su et al. (2024); Dai et al. (2024), we conduct tests on a large number of downstream benchmarks, including zero-shot for BoolQ Clark et al. (2019), HellaSwag Zellers et al. (2019), LAMBADA Paperno et al. (2016), PIQA Bisk et al. (2020), SIQA Sap et al. (2019), StoryCloze Mostafazadeh et al. (2016), and Arc-e Bhakthavatsalam et al. (2021), as well as 5-shot for TriviaQA Joshi et al. (2017) and WebQs Berant et al. (2013). Among them, TriviaQA and WebQs use exact match as the metric, while the remaining benchmarks are evaluated based on accuracy. For a fair comparison, we use the open-source evaluation tool lm-evaluation-harness [^2] https://github.com/EleutherAI/lm-evaluation-harness for assessment.
@@ 4.4.1 Perplexity Results
Table 1 shows the main results of language modeling on the Pile validation sets. With the same parameter activations maintained, MoE models consistently achieve improvements compared to the dense model. In comparison to SMoE and Hash Layer, under a single MoE layer setup, MaskMoE has reduced the PPL on the Pile Validation by 0.14 and 0.08 respectively. In a 12-MoE layers setup, the PPL reductions are 0.07 and 0.05 respectively. Compared to Share-MoE, MaskMoE decreases by 0.24 in single MoE layers and by 0.04 in 12-MoE layers. These results demonstrate the rationality of the MaskMoE design.
It is noteworthy that Share-MoE achieves worse results than SMoE and Hash Layer in the single MoE setup. However, MaskMoE consistently maintains superior performance, whether in single or multi-MoE layers. This indicates that MaskMoE exhibits stronger robustness across various settings.
@@ 4.4.2 Benchmark Results
As shown in Table 2, we present the performance of the models on downstream tasks. Firstly, we observe that MoE models significantly outperform Dense models in downstream tasks. Then, we find that MaskMoE substantially outperforms the other MoE baselines, whether in single-layer or 12-layers settings.
Compared to SMoE, MaskMoE outperforms in all 9 benchmarks with a single MoE layer and in 8 out of 9 benchmarks with 12-MoE layers. Against Hash Layer, MaskMoE excels in 7 out of 9 benchmarks with a single MoE layer and in 8 out of 9 with 12-MoE layers. Compared to Share-MoE, MaskMoE shows significant improvements in 7 out of 9 benchmarks with a single MoE layer and in 6 out of 9 with 12-MoE layers.
This indicates that the performance enhancement is achieved by using our proposed router mask to control the visibility of experts for different tokens. We believe the improvements are due to two factors. Firstly, compared to SMoE and Share-MoE, infrequent tokens are only routed to a single expert, ensuring more thorough training for these tokens. Secondly, in contrast to the Hash Layer, frequent tokens have multiple expert options, which promotes diversity in the representation of frequent tokens.
@@ 5 Analyses
We conduct further experiments to provide more insightful analyses on the proposed MaskMoE. Considering the constraints of computational resources, following previous works Roller et al. (2021); Xie et al. (2023b); Dai et al. (2022), unless otherwise specified, most of the analytical experiments are conducted under the setting of a single MoE layer.
@@ 5.1 Impact of Router Mask
Table 3: Impact of Router Mask. $\spadesuit$ denotes SMoE, $\clubsuit$ denotes Hash Layer. To make the differences in MaskMoE under various configurations more apparent, the values shown in the table are precise to three decimal places.
To visually demonstrate the impact of the routing masking mechanism, the experiments in this subsection do not employ a shared expert structure. Instead, they adopt the same structure as SMoE and Hash Layer, with each MoE layer comprising 64 experts, and the FNN layer is full-sized.
The performance of MaskMoE is predominantly influenced by two hyperparameters: the boundary threshold $P$ for the division of frequent and infrequent tokens, and the maximum number of visible experts $V$ for tokens. As shown in Table 3, we conduct an experimental parameter search for both and reported the Pile Validation PPL. $V_{a}$ represents the number of experts visible for frequent tokens, and $V_{b}$ represents the number of experts visible for infrequent tokens. Notably, when $V_{a}=V_{b}=1$, MaskMoE is equivalent to the Hash Layer, and when $V_{a}=V_{b}=64$, MaskMoE is equivalent to SMoE.
We have the following observations. Firstly, we find that MaskMoE outperforms the Hash Layer. With the $V_{a}=8$ and $V_{b}=1$ settings, MaskMoE reduces perplexity by 0.052 on the Pile validation set compared to the Hash Layer, i.e., $V_{a}=V_{b}=1$. This indicates that MaskMoE’s use of multiple expert encodings for frequent tokens enhances representation diversity, which is beneficial for model performance.
Secondly, compared to SMoE, concentrating tokens into fewer visible experts, such as settings like $V_{a}=V_{b}=8$ or $V_{a}=V_{b}=4$, yields better outcomes. This indicates that routing tokens more densely to a reduced number of visible experts enhances thorough training and improves model performance. Moreover, we observe that the configuration $V_{a}=8$ and $V_{b}=1$ outperforms $V_{a}=V_{b}=8$. This indicates that further reducing the number of visible experts is more beneficial for infrequent tokens.
Lastly, it is noteworthy that for MaskMoE, when $V_{b}=1$, regardless of the value of $V_{a}$ (whether it is 4, 8, 16, or 64), the performance of MaskMoE consistently outperforms that of SMoE and Hash Layer. This demonstrates the robustness of MaskMoE.
5.2 Impact of Shared Expert
To explore the impact of shared experts, in this subsection, we conduct ablation experiments on shared experts, and the results are shown in Table 4. We separately report the average scores of the benchmarks and the PPL of Pile Validation. It is noteworthy that after removing the shared experts, we adopt the same architecture as SMoE and Hash Layer, where the MoE layer consists of 64 fully-sized FFN.
We find that even without shared experts, MaskMoE outperforms SMoE, Hash Layer, and Share-MoE in average benchmark scores and Pile Validation PPL. This further validates the soundness of MaskMoE’s design. Additionally, we observe that incorporating shared expert structures continuously improves MaskMoE’s performance. We believe this is because the shared expert structure promotes greater specialization in token representation DeepSeek-AI et al. (2024).
Model Avg($\uparrow$) PPL($\downarrow$)
SMoE 43.87 6.62
Hash Layer 44.56 6.56
Share-MoE 44.25 6.72
MaskMoE 44.99 6.48
w/o Shared Experts 44.85 6.51
Table 4: Average scores of benchmarks and perplexity on the Pile validation. The best score is marked in bold, and the second best is underlined.
@@ 5.3 Number of Experts
(a) Non-Shared Expert Structures
(b) Shared Expert Structures
@@
To investigate the variation in model performance under different configurations of expert numbers, we conducted experiments with 16, 64, and 128 experts. The experimental results, as shown in Figure 2, reveal the following observations.
Firstly, we observe that for SMoE and Share-MoE, increasing the number of experts does not necessarily lead to a lower perplexity. In fact, when the number of experts reaches 128, the model exhibits a higher PPL compared to having only 64 experts. We believe this is because with more experts, the average number of tokens assigned to each expert decreases, leading to further underfitting of the model. In contrast, the performance of the Hash Layer and MaskMoE consistently improves with the increase in the number of experts, indicating that they are more favorable for scaling the number of experts.
Secondly, we observe that compared to the Hash Layer, MaskMoE consistently performs better across all configurations. Although the Hash Layer also enables experts to be adequately trained with the tokens assigned to them, MaskMoE excels in the diversity of representations for frequent tokens. This superiority in handling token diversity contributes to MaskMoE’s overall advantage over Hash Layer.
Lastly, we find that regardless of the number of experts, MaskMoE consistently maintains the best performance. This indicates that MaskMoE has strong parameter robustness.
@@ 6 Conclusions
In this paper, we analyze dynamic routing models such as SMoE, which routes the same token to multiple experts, potentially causing the model to underfit, especially for infrequent tokens. While fixed routing can mitigate this issue, it struggles with capturing rich features for frequent tokens. To address these issues, we propose a routing masking method aimed at ensuring adequate learning of infrequent tokens while maintaining rich representations for frequent tokens. Extensive experiments demonstrate that the proposed MaskMoE outperforms various downstream tasks and significantly improves perplexity (PPL) on the Pile dataset compared to SMoE, Hash Layer, and Share-MoE.

### v2

@@ 3.3 Load Balance Loss
MoE models typically require distributed training, where different experts are deployed across various computation nodes (e.g., GPU server). Distributed training can lead to load imbalance, where a minority of expert nodes handle the majority of tokens while others remain idle most of the time, and it can negatively impact training efficiency. It is generally desirable for the number of tokens processed by different experts to be roughly equal. To achieve that, a load balancing loss is commonly introduced in the training of MoE models. We follow previous works Huang et al. (2024); Fedus et al. (2022) and adopt a widely used loss as follows.
@@
Where $T$ represents the number of tokens in a mini-batch. $\mathbf{p}^{j}$ denotes the probability distribution of the routing output for the $j$-th token, while $\mathbf{p}_{i}^{j}$ represents the specific probability value of the $i$-th expert. It is noteworthy that for infrequent tokens, which have only one visible expert, the load balancing loss does not apply to them. Since infrequent tokens are assigned to experts through a uniform random hashing, the overall load is supposed to be natively balanced, thus having no need for the balancing loss. The balancing loss primarily regulates the routing of frequent tokens.
@@ 4 Experiments
Table 1: Perplexity (PPL) results of language modeling.
@@ 4.1 Pre-training Dataset
Following previous works, we use the Pile dataset Gao et al. (2021) as pre-training data. The Pile dataset is a large-scale, publicly available corpus, containing 22 domains and over 825GB of English text data. For our experiments, we use the well-known LLaMA tokenizer for tokenization, with a vocabulary size of $32k$. We follow  Xie et al. (2023a); Su et al. (2024) to calculate the sampling rate for each domain based on the number of tokens after tokenization. Due to the computational budget, we follow the pretraining settings of  Xie et al. (2023a); Su et al. (2024); Huang et al. (2024); Xiong et al. (2024); Lian et al. (2024), all models are pre-trained with 100B tokens.
Then, to identify infrequent tokens and frequent tokens, we calculate the frequency of each token in the training set and sort them in descending order of frequency. We categorize the top tokens that cover $P\times 100\%$ of the dataset as frequent ones, and the remaining $(1-P)\times 100\%$ as infrequent ones, where $P$ is a tunable hyperparameter, ranging from 0 to 1.
4.2 Compared Models
We compare the proposed MaskMoE with four models for validation experiments.
  • Dense represents a standard Transformer language model.
  • Hash Layer Roller et al. (2021) signifies that through a random hash method, each token is assigned a fixed expert prior to training. Unless otherwise specified, each MoE layer has 64 experts.
  • MaskMoE aims to achieve a balance between representation diversity in MoE layers and thorough parameter training, via the proposed routing masking method. Similarly to Share-MoE, MaskMoE also employs the remarkable shared expert architecture, which has been proven by  Dai et al. (2024); Rajbhandari et al. (2022) to be beneficial for the training of MoE models.
@@ 4.3 Experimental Setup
Following  Touvron et al. (2023a, b); Huang et al. (2024), we adopt the LLaMA architecture with 24 transformer layers and a hidden layer dimension of 1024 and we employ the AdamW  Loshchilov and Hutter (2019) optimizer with a cosine learning rate decay schedule.
For the MoE model, following  Roller et al. (2021), we conduct experiments under both single-layer and multi-layer settings. For the single-layer setting, we replace the final FFN sublayer with the MoE sublayer. For the multi-layer MoE models, we follow Gshard Lepikhin et al. (2021), and the FFN sublayer with MoE sublayer for MoE layers every other layer, resulting in a total of $12$ MoE layers in this setting. For single-layer MoE models and dense models, we use a learning rate of $3e^{-4}$, while for $12$-layer MoE models, we employ a learning rate of $1e^{-4}$ to ensure stable convergence of the models.
If not specified otherwise, tokens that account for the top 40% of frequency are considered frequent tokens, while the remaining are considered infrequent tokens, i.e., $P=0.4$. Empirically, for frequent tokens, the visible expert count $V$ is set to 8, and for infrequent tokens, the visible expert count $V$ is set to 1. All of our implementations are based on the DeepSpeed [^1] https://github.com/microsoft/DeepSpeed library Rajbhandari et al. (2022); Rasley et al. (2020), which offers robust support for MoE-distributed training. By default, we enable the random token selection method Kim et al. (2021) implemented in the library to facilitate faster model convergence and better runtime efficiency. Following previous works Roller et al. (2021); Fedus et al. (2022); Dai et al. (2022), unless otherwise specified, our experiments only select the top expert.
@@
Table 2: Performances of language models on downstream tasks. The best score is marked in bold, and the second best is underlined.
@@ 4.4 Main Results
In this section, we first present the model’s PPL on the Pile validation set. Then, following Touvron et al. (2023a); Brown et al. (2020); Su et al. (2024); Dai et al. (2024), we conduct tests on a large number of downstream benchmarks, including zero-shot tests for BoolQ Clark et al. (2019), HellaSwag Zellers et al. (2019), LAMBADA Paperno et al. (2016), PIQA Bisk et al. (2020), SIQA Sap et al. (2019), StoryCloze Mostafazadeh et al. (2016), and Arc-e Bhakthavatsalam et al. (2021), as well as 5-shot tests for TriviaQA Joshi et al. (2017) and WebQs Berant et al. (2013). Among them, TriviaQA and WebQs use exact match as the metric, while the remaining benchmarks are evaluated based on accuracy. For a fair comparison, we use the open-source evaluation tool lm-evaluation-harness [^2] https://github.com/EleutherAI/lm-evaluation-harness for assessment.
@@ 4.4.1 Perplexity Results
Table 1 shows the main results of language modeling on the Pile validation sets. With the same number of activated parameters during inference (i.e., 468M), MoE models consistently achieve improvements over the dense model. In comparison to SMoE and Hash Layer, under the Single-MoE-Layer setup, MaskMoE has respectively reduced the PPL on the Pile Validation by 0.14 and 0.08. In the Multi-MoE-Layer setup, their corresponding PPL are, respectively, reduced by 0.07 and 0.05. Compared to Share-MoE, MaskMoE achieves a PPL decrease of 0.24 in Single-MoE-Layer setup and another PPL decrease of 0.04 in Multi-MoE-Layer setup. Such results demonstrate the reasonableness of the MaskMoE design.
It is noteworthy that Share-MoE achieves worse results than SMoE and Hash Layer in the Single-MoE-Layer setup. However, MaskMoE consistently maintains superior performance, whether in single or multi-MoE layers. It indicates that MaskMoE exhibits stronger robustness across various settings.
@@ 4.4.2 Benchmark Results
As shown in Table 2, we present the performance of the models on downstream tasks. Firstly, we observe that MoE models consistently and significantly outperform the dense models in downstream tasks. And more importantly, we find that MaskMoE substantially outperforms the other MoE baselines, whether in Single-MoE-Layer setup or Multi-MoE-Layer setup.
Compared to SMoE, MaskMoE consistently outperforms it in all 9 benchmarks with the Single-MoE-Layer setup and 8 out of 9 benchmarks with the Multi-MoE-Layer setup. Against Hash Layer, MaskMoE excels in 7 out of 9 benchmarks with the Single-MoE-Layer setup and in 8 out of 9 benchmarks with the Multi-MoE-Layer setup. Compared to Share-MoE, MaskMoE demonstrates significant improvements in 7 out of 9 benchmarks with the Single-MoE-Layer setup and in 6 out of 9 benchmarks with the Multi-MoE-Layer setup.
The result above well demonstrates the performance enhancement achieved by using our proposed router masking method to control the visibility of experts for different tokens. We believe the improvements are due to two factors. Firstly, compared to SMoE and Share-MoE, infrequent tokens are only routed to a single expert, ensuring more thorough training for those tokens. Secondly, in contrast to the Hash Layer, frequent tokens have multiple expert options, which promotes diversity in the representation learning of frequent tokens.
@@ 5 Analyses
We conduct further experiments to provide more insightful analyses on the proposed MaskMoE. Considering the constraints of computational resources and following previous works Roller et al. (2021); Xie et al. (2023b); Dai et al. (2022), unless otherwise specified, most of the analytical experiments are conducted under the setting of a single MoE layer.
5.1 Impact of Shared Expert
To explore the impact of shared experts, we conduct ablation experiments on them (i.e., removing the shared experts), and the results are shown in Table 3. We separately report the average scores [^3] Due to the limited space, here we simply average all metrics on downstream tasks to get an average score for each model. of the benchmarks and the PPL of Pile Validation. Note that after removing the shared experts, we adopt the same architecture as SMoE and Hash Layer, where the MoE layer consists of 64 fully-sized FFN.
We find that even without shared experts, MaskMoE still outperforms SMoE, Hash Layer, and Share-MoE in average benchmark scores and PPL. Such a result further validates the soundness of our proposed MaskMoE. Moreover, incorporating shared expert structures continuously improves MaskMoE’s performance. We attribute it to that the shared expert structure promotes greater specialization in token representation DeepSeek-AI et al. (2024).
@@ 5.2 Impact of Router Mask
Model Avg($\uparrow$) PPL($\downarrow$)
SMoE 43.87 6.62
Hash Layer 44.56 6.56
Share-MoE 44.25 6.72
MaskMoE 44.99 6.48
w/o Shared Experts 44.85 6.51
Table 3: Average scores benchmarks and PPL on the Pile validation sets. The best score is marked in bold, and the second best is underlined.
Considering the proposed MaskMoE without shared experts still outperforms the compared MoE methods, to show the impact of router mask more significantly and minimize the impact of the shared expert, the experiments remove the shared expert. Instead, we adopt the same structure as SMoE and Hash Layer, with each MoE layer comprising 64 full-sized FNN modules as experts.
The performance of MaskMoE is predominantly influenced by two hyperparameters: the boundary threshold for the separation of frequent and infrequent tokens, and the maximum number of visible experts (i.e., $V$) for tokens. As shown in Table 4, we conduct an experimental parameter search for both and report their PPL values on the Pile validation sets. $V_{a}$ represents the number of experts visible for frequent tokens, and $V_{b}$ represents the number of experts visible for infrequent tokens. Notably, when $V_{a}=V_{b}=1$, MaskMoE is equivalent to the Hash Layer, and when $V_{a}=V_{b}=64$, MaskMoE is equivalent to SMoE.
Then we have the following observations. Firstly, we find that MaskMoE outperforms the Hash Layer. With $V_{a}=8$ and $V_{b}=1$, MaskMoE reduces ppl by 0.052 on the Pile validation set compared to the Hash Layer, i.e., $V_{a}=V_{b}=1$. The result indicates that the multiple-expert design in the proposed MaskMoE for frequent tokens does enhance their representation diversity, which is beneficial for model performance.
Secondly, compared to SMoE, concentrating tokens into fewer visible experts, such as settings like $V_{a}=V_{b}=8$ or $V_{a}=V_{b}=4$, yields better outcomes. It indicates that routing tokens more densely to a reduced number of visible experts can enhance thorough training and improve model performance. Moreover, we observe that the configuration $V_{a}=8$ and $V_{b}=1$ outperforms $V_{a}=V_{b}=8$. It well verifies that further reducing the number of visible experts is more beneficial for infrequent tokens.
Finally, it is noteworthy that for MaskMoE, when $V_{b}=1$, regardless of the value of $V_{a}$ (whether it is 4, 8, 16, or 64), the performance of MaskMoE consistently outperforms that of SMoE and Hash Layer. That also demonstrates the robustness of MaskMoE.
@@
Table 4: Impact of Router Mask. $\spadesuit$ denotes SMoE, $\clubsuit$ denotes Hash Layer.
@@ 5.3 Number of Experts
To investigate the variation in model performance under different configurations of expert numbers, we conducted experiments with 32, 64, and 128 experts. The experimental results, as shown in Figure 2, reveal the following observations.
First of all, we observe that for SMoE and Share-MoE, increasing the number of experts does not necessarily lead to a lower PPL. In fact, when the number of experts reaches 128, both models even exhibit higher PPLs compared to having only 64 experts. We attribute it to that having more experts leads to the occurrences of an identical token being spread among more experts, resulting in each expert not learning enough about this token, which leads to model underfitting. In contrast, the performance of Hash Layer and MaskMoE continue to improve with the increase in the number of experts. We attribute it to that, despite the growing number of experts, the visible experts for an identical token in Hash Layer and MaskMoE remain a fixed, small number, allowing the experts to continue learning adequately. It indicates that they are more favorable for scaling the number of experts.
Secondly, we observe that compared to the Hash Layer, MaskMoE consistently performs better across all configurations. Although the Hash Layer also enables experts to be adequately trained with the tokens assigned to them, MaskMoE excels in the diversity of representations learning for frequent tokens. such a superiority in handling token diversity contributes to MaskMoE’s overall advantage over Hash Layer.
Finally, we find that regardless of the number of experts, MaskMoE consistently maintains the best performance. That further verifies the superiority of the proposed MaskMoE and the reasonableness of its design.
(a) None-Shared Expert Structures
(b) Shared-Expert Structures
@@ 6 Conclusions
In existing MoE language models, the widely-used dynamic routing models such as SMoE, which routes the same token to multiple experts, potentially cause the model to encounter underfitting during training, especially for infrequent tokens. Although fixed routing like Hash Layer can mitigate the routing fluctuation issue, it struggles with capturing rich features for frequent tokens. To tackle the routine fluctuation issue and the representation diversity learning issue at the meantime, we propose a routing masking method termed MaskMoE, aiming at ensuring adequate learning of infrequent tokens while maintaining rich representations learning for frequent tokens. Extensive experiments demonstrate that the proposed MaskMoE outperforms existing MoE methods like SMoE, Hash Layer, and Share-MoE on various downstream tasks and significantly reduces PPL on the Pile validation sets. That will demonstrates the effectiveness of MaskMoE.